# NLO corrections and factorization for transverse single-spin asymmetries

### v1

@@ Abstract
In the present paper, we will address the SSA for the single-inclusive process $\ell p^{\uparrow}\to hX$, with an incoming transversely polarized proton and a hadron $h$ produced at large transverse momentum $P_{T,h}$. Note that we assume the scattered lepton to remain undetected. At sufficiently high energy, the hadron may also be replaced by an observed hadronic jet. The asymmetry is defined as
@@
where $\bm{S}_{T}$ denotes the transverse proton spin vector and $(E_{h},\bm{P}_{h})$ the hadron’s four-momentum. The unexpectedly large values of $A_{UT}$ found in measurements for the single-inclusive reaction $pp^{\uparrow}\to hX$ [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] in fact marked the beginning of this field.
A key property of single-inclusive reactions is that they are characterized by a single large momentum scale, here $P_{T,h}$. Therefore, collinear factorization is expected to provide the appropriate framework for the analysis. This should also apply to the spin-dependent cross section in the numerator of $A_{UT}$, even though it is power-suppressed as $1/P_{T,h}$ relative to the unpolarized one. Indeed, it was argued early on [13, 14, 15, 16, 17, 18] that $\Delta\sigma$ may be factorized in terms of universal collinear “twist-3” hadronic matrix elements convoluted with partonic hard-scattering functions that are computable in QCD perturbation theory and are specific to the SSA under consideration. The hadronic matrix elements appear as correlation functions of partons in the proton, but also in parton fragmentation to the observed hadron.
@@
×$\times\langle P,S|\,\bar{q}_{j}(0)\,[0,\mu n]\,ig\,G^{n\rho}(\mu n)\,[\mu n,\lambda n]\,q_{i}(\lambda n)\,|P,S\rangle.$
It consists of “dynamical” matrix elements of two quark field components $q_{i}$, $q_{j}$ and a gluonic field strength tensor $G$, separated by Wilson lines denoted by []delimited-[][$\ldots$][ … ]. As indicated, the correlator depends on two light-cone momentum fractions $x$ and $x^{\prime}$. As a full matrix in Dirac space, it can be parameterized as [27]
@@
The support of the functions $F_{(qg)}^{\mathbf{1,5}}(x,x^{\prime})$ is given by the conditions $-1\leq x,x^{\prime}\leq 1$ and $|x-x^{\prime}|\leq 1$. Figure 1 shows the ensuing region in the $x,x^{\prime}$ plane. In general, the functions will enter a factorized cross section with an integration over all kinematically accessible $x$ and $x^{\prime}$. However, as is well known, a non-vanishing single-spin asymmetry requires the presence of a phase in the scattering process. Such a phase may be generated when a propagator (momentum $q$) in the partonic hard-scattering goes on-shell, via the relation
@@
  • There are integrations over both $x$ and $x^{\prime}$. Such contributions arise when $q$ in Eq. (4) contains the momentum of an unobserved final-state parton which has to be integrated over via a phase space integral. We refer to them as integral (Int) contributions.
@@
×$\times\sum_{\mathcal{P}}\sum_{(ij),k}\sum_{\mathbf{I}=\mathbf{1,5}}\hat{\sigma}_{(ij)\to k}^{\mathcal{P},\mathbf{I}}(v,w,\zeta)\,{\cal F}_{(ij)}^{\mathcal{P},\mathbf{I}}(x,\zeta x)\,D_{k}^{h}(z),\;\;\;$
@@
The central result of this paper is that upon combination of all terms we arrive at finite results for all partonic channels contributing at NLO to the spin-dependent cross section. This validates factorization at collinear twist-3 to order $\alpha_{s}$. For single-inclusive jet production, our result presents the complete NLO answer; we recall that for hadron production there will also be twist-3 fragmentation contributions that we do not address here [29, 27].
The final expressions for the NLO corrections are lengthy and will be reported in a companion paper. Here we only present the leading double-logarithmic threshold contribution for $w\to 1$, which is entirely SGP and reads:
@@
Phenomenology of single-spin asymmetry at NLO. We now present a few exploratory numerical results for the single-spin asymmetry in $ep^{\uparrow}\to\pi^{\pm}X$, adopting kinematics relevant for comparison to HERMES data [34] taken at c.m.s. energy $\sqrt{s}=7.25$ GeV. For the spin-dependent cross section we need input distributions for the quark-gluon-quark correlation functions $F_{(qg)}^{\mathbf{1,5}}$ for all flavors $q$, as well as for the three-gluon correlation functions. Since none of these functions are currently known, a realistic prediction is not feasible. The only constraints are provided by the connection of the Qiu-Sterman functions to the first $k_{\perp}$ moments of the Sivers functions [35], $f_{1T}^{\perp(1),q}(x)=\pi F_{(qg)}^{\mathbf{1}}(x,x)$, along with $F_{(qg)}^{\mathbf{5}}(x,x)=0$, and lattice-QCD computations [36, 37, 38, 39] of the quantity $\int_{-1}^{1}\mathrm{d}x\int_{x-1}^{1}\mathrm{d}x^{\prime}F_{(qg)}^{\mathbf{1}}(x,x^{\prime})$. For our numerical studies we resort to a set of three simple models for $F_{(qg)}^{\mathbf{1,5}}(x,x^{\prime})$ for up and down quarks, assuming all other multiparton correlation functions to vanish. Technically, we write the momentum fractions in polar coordinates, $x=r\cos(\varphi+\tfrac{\pi}{4})$, $x^{\prime}=r\sin(\varphi+\tfrac{\pi}{4})$, and then expand the $\varphi$ dependence of $F_{(qg)}^{\mathbf{1,5}}$ in terms of a Fourier series, selecting three sets of values for the low Fourier modes. Figure 3 shows the three resulting $u$- and $d$-quark functions $F_{(qg)}^{\mathbf{1}}$ as contour plots in the $x,x^{\prime}$ plane. Further details as well as a related plot of $F_{(qg)}^{\mathbf{5}}$ are given in our companion paper. We stress that each one of the models in Fig. 3 respects the Sivers function and lattice constraints described above. To compute the $k_{\perp}$ moments of the Sivers functions we adopt the analysis of Ref. [40]. Note that for this exploratory study we do not take into account scale evolution of the twist-3 correlation functions but rather set the scale for our model by the average pion transverse momentum $\langle P_{T,\pi}\rangle\simeq 1.15$ GeV accessed by HERMES. We use the pion fragmentation functions of Ref. [41]. We recall that we do not take into account any twist-3 fragmentation contributions. The unpolarized cross section in the denominator of the asymmetry is computed at NLO, using the results of [31] and the MSTW2008 parameterization [42] for the unpolarized proton’s parton distributions.
@@
Figure 4 shows the asymmetries $A_{UT}$ for $\pi^{+}$ (left) and $\pi^{-}$ (right) production, as functions of the pion’s Feynman variable $x_{F}$. We present the LO result as well as the NLO ones for our three scenarios. Our main finding is that the NLO asymmetry is highly sensitive to the set of twist-3 matrix elements chosen. We found that this is primarily due to an intricate interplay among the various pole contributions. We note that – arguably accidentally – our scenario 1 actually describes the HERMES $\pi^{+}$ data rather well, while none of the scenarios does a particularly good job for $\pi^{-}$. We stress how crucial our NLO corrections are in this context, and how complex the twist-3 framework is: by construction, the LO result – which only probes the diagonal $x=x^{\prime}$ – is the same for each of the three model sets of twist-3 matrix elements. For further comparison, we also show the asymmetry computed with NLO corrections only in the denominator (dashed lines). Thanks to sizable positive NLO corrections for the unpolarized cross section it is strongly reduced relative to LO, as was first observed in [43]. Our results demonstrate that inclusion of NLO corrections also in the numerator of $A_{UT}$ is crucial for phenomenology.
@@
Conclusions. We have presented the NLO QCD corrections to the spin-dependent cross section for the single-inclusive process $\ell p^{\uparrow}\to hX$ with transversely polarized protons. We have demonstrated that collinear twist-3 factorization holds at the one-loop level for the contributions associated with twist-3 nucleon matrix elements. This is a highly non-trivial result, given the complexity of the twist-3 framework for polarized single-inclusive cross sections. Our calculation thus serves as a proof of principle that such NLO twist-3 calculations are possible, including phenomenological evaluations. The NLO single-spin asymmetry $A_{UT}$ shows great sensitivity to the set of twist-3 matrix elements adopted, which also highlights the impact that future precision EIC data might have. Our work presented here opens the door to future full calculations of $pp^{\uparrow}\to hX$.

### v2

@@ Abstract
In the present paper, we will address the SSA for the single-inclusive process $\ell p^{\uparrow}\to hX$, with an incoming transversely polarized proton and a hadron $h$ produced at large transverse momentum $P_{h,T}$. Note that we assume the scattered lepton to remain undetected. At sufficiently high energy, the hadron may also be replaced by an observed hadronic jet. The asymmetry is defined as
@@
where $\bm{S}_{T}$ denotes the transverse proton spin vector and $(E_{h},\bm{P}_{h})$ the hadron’s four-momentum. The unexpectedly large values of $A_{UT}$ found in measurements for the single-inclusive reaction $pp^{\uparrow}\to hX$ [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] in fact marked the beginning of this field.  [^1] The conceptual and kinematical similarity of the asymmetry (1) to the corresponding one in $pp^{\uparrow}\to hX$ is the main motivation for us to study the single-inclusive process $\ell p^{\uparrow}\to hX$ instead of the more conventional semi-inclusive process $\ell p^{\uparrow}\to\ell hX$ (SIDIS).
A key property of single-inclusive reactions is that they are characterized by a single large momentum scale, here $P_{h,T}$. Therefore, collinear factorization is expected to provide the appropriate framework for the analysis. This should also apply to the spin-dependent cross section in the numerator of $A_{UT}$, even though it is power-suppressed as $1/P_{h,T}$ relative to the unpolarized one. Indeed, it was argued early on [13, 14, 15, 16, 17, 18] that $\Delta\sigma$ may be factorized in terms of universal collinear “twist-3” hadronic matrix elements convoluted with partonic hard-scattering functions that are computable in QCD perturbation theory and are specific to the SSA under consideration. The hadronic matrix elements appear as correlation functions of partons in the proton, but also in parton fragmentation to the observed hadron.
@@
$\hskip-48.36958pt\times\langle P,S|\,\bar{q}_{j}(0)\,[0,\mu n]\,ig\,G^{n\rho}(\mu n)\,[\mu n,\lambda n]\,q_{i}(\lambda n)\,|P,S\rangle.$
It consists of “dynamical” matrix elements of two quark field components $q_{i}$, $q_{j}$ and a gluonic field strength tensor $G$, separated by Wilson lines denoted by $\ldots$]. As indicated, the correlator depends on two light-cone momentum fractions $x$ and $x^{\prime}$. As a full matrix in Dirac space, it can be parameterized as [27]
@@
The support of the functions $F_{(qg)}^{\mathbf{1,5}}(x,x^{\prime})$ is given by the conditions $1\leq x,x^{\prime}\leq 1$ and $|x-x^{\prime}|\leq 1$. Figure 1 shows the ensuing region in the $x,x^{\prime}$ plane. In general, the functions will enter a factorized cross section with an integration over all kinematically accessible $x$ and $x^{\prime}$. However, as is well known, a non-vanishing single-spin asymmetry requires the presence of a phase in the scattering process. Such a phase may be generated when a propagator (momentum $q$) in the partonic hard-scattering goes on-shell, via the relation
@@
  • There are integrations over both $x$ and $x^{\prime}$. Such contributions arise when $q$ in Eq. (4) contains the momentum of an unobserved final-state parton which has to be integrated over via a phase space integral. We refer to them as integral (Int) contributions  [^2] We note that for these contributions typically direct application of (4) results in $q$ appearing in the arguments of the functions $F_{(qg)}^{\mathbf{1,5}}(x,x^{\prime})$, which clearly is not desirable. It is easier in such cases to first carry out the integration over the momentum of the unobserved particle and take the imaginary part afterwards..
@@
$\hskip-45.52458pt\times\sum_{\mathcal{P}}\sum_{(ij),k}\sum_{\mathbf{I}=\mathbf{1,5}}\hat{\sigma}_{(ij)\to k}^{\mathcal{P},\mathbf{I}}(v,w,\zeta)\,{\cal F}_{(ij)}^{\mathcal{P},\mathbf{I}}(x,\zeta x)\,D_{k}^{h}(z),\;\;\;$
@@
The central result of this paper is that upon combination of all terms we arrive at finite results for all partonic channels contributing at NLO to the spin-dependent cross section [33].This validates factorization at collinear twist-3 to order $\alpha_{s}$. For single-inclusive jet production, our result presents the complete NLO answer; we recall that for hadron production there will also be twist-3 fragmentation contributions that we do not address here [29, 27].
The final expressions for the NLO corrections are lengthy and will be reported in a companion paper [34]. Here we only present the leading double-logarithmic threshold contribution for $w\to 1$, which is entirely SGP and reads:
@@
Phenomenology of single-spin asymmetry at NLO. We now present a few exploratory numerical results for the single-spin asymmetry in $ep^{\uparrow}\to\pi^{\pm}X$, adopting kinematics relevant for comparison to HERMES data [36] taken at c.m.s. energy $\sqrt{s}=7.25$ GeV. For the spin-dependent cross section we need input distributions for the quark-gluon-quark correlation functions $F_{(qg)}^{\mathbf{1,5}}$ for all flavors $q$, as well as for the three-gluon correlation functions. Since none of these functions are currently known, a realistic prediction is not feasible. The only constraints are provided by the connection of the Qiu-Sterman functions to the first $k_{\perp}$ moments of the Sivers functions [37], $f_{1T}^{\perp(1),q}(x)=\pi F_{(qg)}^{\mathbf{1}}(x,x)$, along with $F_{(qg)}^{\mathbf{5}}(x,x)=0$, and lattice-QCD computations [38, 39, 40, 41] of the quantity $\int_{-1}^{1}\mathrm{d}x\int_{x-1}^{1}\mathrm{d}x^{\prime}F_{(qg)}^{\mathbf{1}}(x,x^{\prime})$. For our numerical studies we resort to a set of three simple models for $F_{(qg)}^{\mathbf{1,5}}(x,x^{\prime})$ for up and down quarks, assuming all other multiparton correlation functions to vanish. Technically, we write the momentum fractions in polar coordinates, $x=r\cos(\varphi+\tfrac{\pi}{4})$, $x^{\prime}=r\sin(\varphi+\tfrac{\pi}{4})$, and then expand the $\varphi$ dependence of $F_{(qg)}^{\mathbf{1,5}}$ in terms of a Fourier series, selecting three sets of values for the low Fourier modes. Figure 3 shows the three resulting $u$- and $d$-quark functions $F_{(qg)}^{\mathbf{1}}$ as contour plots in the $x,x^{\prime}$ plane. Further details as well as a related plot of $F_{(qg)}^{\mathbf{5}}$ are given in our companion paper [34]. We stress that each one of the models in Fig. 3 respects the Sivers function and lattice constraints described above. To compute the $k_{\perp}$ moments of the Sivers functions we adopt the analysis of Ref. [42]. Note that for this exploratory study we do not take into account the proper scale evolution of the twist-3 correlation functions but only vary the scale of the unpolarized $f_{1}^{q}$ parton distributions underlying our model. We then generate theoretical uncertainty bands by varying the scale in the range $1~{\mathrm{GeV}}\leq\mu\leq 2.2~{\mathrm{GeV}}$ corresponding to the boundaries of the HERMES $P_{\pi,T}$ bin, using $\mu=\langle P_{\pi,T}\rangle\simeq 1.15$ GeV as the central value. We use the pion fragmentation functions of Ref. [43]. We recall that we do not take into account any twist-3 fragmentation contributions. The unpolarized cross section in the denominator of the asymmetry is computed at NLO, using the results of [31] and the MSTW2008 parameterization [44] for the unpolarized proton’s parton distributions.
@@
Figure 4 shows the asymmetries $A_{UT}$ for $\pi^{+}$ (left) and $\pi^{-}$ (right) production, as functions of the pion’s Feynman variable $x_{F}$. We present the LO result as well as the NLO ones for our three scenarios. Our main finding is that the NLO asymmetry is highly sensitive to the set of twist-3 matrix elements chosen, and the preliminary scale uncertainty bands do not overlap over a wide range in $x_{F}$. We found that this is primarily due to an intricate interplay among the various pole contributions. We note that – arguably accidentally – our scenario 1 actually describes the HERMES $\pi^{+}$ data rather well, while none of the scenarios does a particularly good job for $\pi^{-}$. We stress how crucial our NLO corrections are in this context, and how complex the twist-3 framework is: by construction, the LO result – which only probes the diagonal $x=x^{\prime}$ – is the same for each of the three model sets of twist-3 matrix elements. For further comparison, we also show the asymmetry computed with NLO corrections only in the denominator (dashed lines). Thanks to sizable positive NLO corrections for the unpolarized cross section it is strongly reduced relative to LO, as was first observed in [45]. Our results demonstrate that inclusion of NLO corrections also in the numerator of $A_{UT}$ is crucial for phenomenology.
@@
Conclusions. We have presented the NLO QCD corrections to the spin-dependent cross section for the single-inclusive process $\ell p^{\uparrow}\to hX$ with transversely polarized protons. We have demonstrated that collinear twist-3 factorization holds at the one-loop level for the contributions associated with twist-3 nucleon matrix elements. This is a highly non-trivial result, given the complexity of the twist-3 framework for polarized single-inclusive cross sections, compared to that for the Drell-Yan process and semi-inclusive DIS. In fact, this is the first time that it has been possible to obtain the NLO corrections for a single-inclusive single-spin asymmetry with its all-important derivative terms. Our calculation thus serves as a proof of principle that such NLO twist-3 calculations are possible, including phenomenological evaluations. The NLO single-spin asymmetry $A_{UT}$ shows great sensitivity to the set of twist-3 matrix elements adopted, which also highlights the impact that future precision EIC data might have. We hope that extended data coverage at the EIC will make it possible to better constrain the twist-3 correlation functions also for $x\neq x^{\prime}$ so that one can go beyond just modeling this regime. This, along with full use of evolution of the functions, will remain a challenge for the future. In any case, our work presented here opens the door to future full calculations of $pp^{\uparrow}\to hX$.